OpenStudy (anonymous):

MEDAL GIVEN!! please explain properly Suppose a traveler wanted to visit a museum, an art gallery, and the state capitol building. 45-minute tours are offered at each attraction hourly from 10 a.m. through 3 p.m. (6 different hours). Solve the problem, disregarding travel time. In how many ways could the traveler schedule two of the three tours in one day?

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@superman38

OpenStudy (anonymous):

You have 6 time slots to choose 3 from. which is simply: 6 * 5 * 4 = 120

OpenStudy (anonymous):

wait so the art gallery and the museum has nothing to do with it?

OpenStudy (anonymous):

naw

OpenStudy (anonymous):

The times 10 to 3 are the starting times of the tours, so there are 6 of them. You have 6 time slots to choose 3 from. which is simply: 6 * 5 * 4 = 120

OpenStudy (anonymous):

oh okay, I have one more question

OpenStudy (anonymous):

go ahead lil ma

OpenStudy (anonymous):

In how many ways can the traveler visit all three places in one day?

OpenStudy (anonymous):

each hour a tour starts at each place and the tours are less then an hour. now we figure out the poisble number of things that could be scheduled for each hour with the musem tour being after noon, and not going on a tour a fourth option so for the 10 am hour 3 poisabitiles 11 am hour 3 poisabitiles noon 4 poisabities 1 pm hour 3 poisabitiles 2 pm hour 2 poisabities 3 pm hour 1 poisabitie only for noon is there all 4 optoins (one of the three tours or no tour) and every hour after that there is one less optoin, because the day is coming to an end. next step, I belive is to mutiply the optoins for each hour 3*3*4*3*2*1= 216

OpenStudy (anonymous):

no that's not what is asking

OpenStudy (anonymous):

In how many ways can the traveler visit all three places in one day?

OpenStudy (anonymous):

does it have answer choices

OpenStudy (anonymous):

i wouuld say 3 ways

OpenStudy (anonymous):

no it doesn't have choices

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

In how many ways could the traveler schedule all three tours before 1 p.m.?

OpenStudy (anonymous):

1 way

OpenStudy (anonymous):

the answers says 6

OpenStudy (anonymous):

oh six

OpenStudy (anonymous):

can you explain

OpenStudy (anonymous):

oh damn dats kinda to do so

OpenStudy (anonymous):

can you try to explain

OpenStudy (anonymous):

im not good at tht im only good at givnign answers but not explantions

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

im sorry

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

ok